OpenStudy (anonymous):

A general term of the sequence -2, -1, -5, 11, -53, 203,... can be expressed as (a-b^(n-1))/5 where a and b are both integers. Find the value of ab.

Parth (parthkohli):

\[n = 1 \Rightarrow T_1 = \frac{a - 1}{5} = -2\]

Parth (parthkohli):

When you find out the value of \(a\), you can then plug in \(n = 2\) and find \(b\).

OpenStudy (anonymous):

I get ab = 36.

Parth (parthkohli):

\[-9 - b = -5\]\[b = -4\]Good job!